hi how can i find the answer to a question like 2y - 4x = 8 and then decide whether the given ordered pair (-2,8). i think i have to find if x=? and y= ? and the two ? will b the numbers -2 and 8, but im not sure. canyou help me? thank you!
Answer by checkley77(12844)

2y - 4x = 8
(-2,8)
Replace the x & y terms with -2 & 8 & solve.
2*8-4*-2=8
16-8=8
8=8 ans.
YES (-2,8) are solutions to this equation.